In an R-L-C series circuit, the differential equation for the instantaneous charge q(t) on the capacitor is   2 2   d q dq q L R Et dt dt C . Determine the charge q(t) and current i(t) for a circuit with R  10 ohm, L = 1 henry, C = 2 10 farad and E(t) = 50 10 cos t volts. What is the steady-state current for this circuit?


1
Expert's answer
2021-09-23T17:42:15-0400
"1\dfrac{d^2q}{dt^2}+10\dfrac{dq}{dt}+\dfrac{q}{10^{-2}}=50\cos(10t)"

Homogeneous differential equation


"\dfrac{d^2q}{dt^2}+10\dfrac{dq}{dt}+100q=0"

Corresponding (auxiliary) equation


"r^2+10r+100=0"

"D=(10)^2-4(1)(100)=-300"

"r=\dfrac{-10\pm\sqrt{-300}}{2(1)}=-5\pm 5\sqrt{3}i"

The general solution of the homogeneous differential equation is


"q_h=c_1e^{-5t}\cos(5\sqrt{3}t)+c_2e^{-5t}\sin(5\sqrt{3}t)"

Find the particular solution of the non homogeneous differential equation


"q_p=A\cos(10t)+B\sin(10t)"

"q_p'=-10A\sin(10t)+10B\cos(10t)"

"q_p''=-100A\cos(10t)-100B\sin(10t)"

Substitute


"-100A\cos(10t)-100B\sin(10t)-100A\sin(10t)"

"+100B\cos(10t)+100A\cos(10t)+100B\sin(10t)"

"=50\cos(10t)"

"100B=50"

"A=0"

The particular solution of the non homogeneous differential equation


"q_p=\dfrac{1}{2}\sin(10t)"

The general solution of the given differential equation


"q(t)=c_1e^{-5t}\cos(5\sqrt{3}t)+c_2e^{-5t}\sin(5\sqrt{3}t)+\dfrac{1}{2}\sin(10t)"

Then


"i(t)=\dfrac{dq}{dt}=-5c_1e^{-5t}\cos(5\sqrt{3}t)-5c_2e^{-5t}\sin(5\sqrt{3}t)"

"-5\sqrt{3}c_2e^{-5t}\sin(5\sqrt{3}t)+5\sqrt{3}c_2e^{-5t}\cos(5\sqrt{3}t)"

"+5\cos(10t)"

"t\to\infin"

 The steady-state current for this circuit is


"i_{steady-state}=5\cos(10t)"